OpenStudy (anonymous):

Can this be factored? x^2 + 10x + 25

OpenStudy (anonymous):

what adds to 10 and when multiplied gives you 25?

OpenStudy (anonymous):

(n + n) = 10 & (n * n) = 25

OpenStudy (anonymous):

=x^2+5x+5x+25 =x(x+5)+5(x+5) =(x+5)(x+5) ==(x+5)^2